1. In a system, executing a non-flow process, the work and heat per degree

change of temperature are given by

dW

dT

= 200 W-s/°C and dQ

dT = 160 J/°C

What will be the change of internal energy of the system when its temperature changes

from

T1 = 55°C to T2 = 95°C ?


1
Expert's answer
2021-09-09T11:02:06-0400

In a system, executing a non flow process.


(NB: non flow process is a process in which no mass transfer takes place between system and surroundings)


From first law of thermodynamics,

Q=W+U

differentiating with respect to time,

"\dfrac{dQ}{dT} = \dfrac{dW}{dT} +\dfrac{ dU}{dT}"


given,

"\dfrac{dQ}{dT} = 200 W.s/°C = 200 J/°C"


"\dfrac{dW}{dT} = 160 J/°C"


so,

"\dfrac{dU}{dT} = 200-160= 40J/°C"


change of internal energy,

"U = \dfrac{dU}{dT}× ∆T= 40 J/°C × (95°C-55°C)

= 40J/°C × 40°C
= 1600\ J"